OpenStudy (anonymous):

Need Help, Factoring. How would you manage to factor this? I need it step by step please. 0 = x^2 + 23 - 420

OpenStudy (anonymous):

Oh, I messed up the typing. Its 0 = x^2 + 23x - 420

OpenStudy (trojanpoem):

\[x^{2} +23x - 420 = 0 \] \[ax^{2} +bx + c = 0 \] a = 1, b = 23, c = -420 use the quadratic formula: \[x = \frac{ -b \pm \sqrt{ b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

Thank you!

OpenStudy (trojanpoem):

Any time.